Grade 11General Physics

a scooter going due east at 10 m/s turns right through an angle 90 . if the speed of the scooter remains unchanged in taking this turn, the change in velocity of the scooter is

Peeyush
11 Years agoGrade 11

4 Answers

Peeyush
11 Years ago
Please someone aswer this question as sonn as someone can
 
Shubh Saraswat
9 Years ago
The answer is 10 m/s in south Eastern direction.... It`s a fucking easy question..... This just a 10 grade vectors question😂😂😂😂
Rakhi
9 Years ago
The answer is 14.14 m/s in s-w directionV1=10i^ v2=10j^ resultant=√(10)sq+(10)sq=√200=10√2=10×1.41=14.14 m/s ans
Yasar arafat
7 Years ago
The answer is 14.14 south-west direction.
Reason: Δv=final velocity - initial velocity
Final velocity =10j
Initial velocity =10i
Δv = 10j -10i
|Δv| = √10^2+ (-10)^2
|Δv| = √200
        = 10√2
       = 10(1.414)
|Δv| = 14.14
Then final velocity - initial velocity comes in south west direction.